OpenStudy (anonymous):

Lisa wants to buy a phone card for $5. Company X charges $0.37 for the first minute and $0.08 for each additional minute. Company Y charges $0.52 for the first minute and $0.14 for each additional minute. If Lisa makes just one call, which phone company gives her more minutes for $5? How many more minutes?

OpenStudy (anonymous):

Company Y equation is Y=0.14x+0.52 Company X equation is Y=0.08x+0.37 Then replace Y with 5 and solve. Company Y: 5=0.14x+0.52 4.48=0.14x 32 minutes Company X: 5=0.08x+0.37 4.63=0.08x about 58 minutes 58-32= 26 Answer: Company X: about 26 minutes.

OpenStudy (anonymous):

yay!!! thanks!!!!!